### v1

# Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing

Abstract
Reinforcement Learning (RL) has empowered Multimodal Large Language Models (MLLMs) to achieve superior human preference alignment in Image Quality Assessment (IQA). However, existing RL-based IQA models typically rely on coarse-grained global views, failing to capture subtle local degradations in high-resolution scenarios. While emerging “Thinking with Images” paradigms enable multi-scale visual perception via zoom-in mechanisms, their direct adaptation to IQA induces spurious “cropping-implies-degradation” biases and misinterprets natural depth-of-field as artifacts. To address these challenges, we propose Q-Probe, the first agentic IQA framework designed to scale IQA to high resolution via context-aware probing. First, we construct Vista-Bench, a pioneering benchmark tailored for fine-grained local degradation analysis in high-resolution IQA settings. Furthermore, we propose a three-stage training paradigm that progressively aligns the model with human preferences, while simultaneously eliminating causal bias through a novel context-aware cropping strategy. Extensive experiments demonstrate that Q-Probe achieves state-of-the-art performance in high-resolution settings while maintaining superior efficacy across resolution scales.
Machine Learning, ICML, IQA, Agentic
[…]
figureChallenges in detecting subtle distortions via global perception versus local zooming. (a) Existing MLLMs fail to capture subtle local artifacts. (b) Even when visible via cropping, Semantic Robustness Bias causes models to ignore defects in key semantic areas (e.g., face). (c-d) Naive zooming leads to Logic Collapse, where the model misinterprets natural bokeh as blur (c) or falsely learns that Zooming implies Low Quality (d). (e) Q-Probe mimics human active viewing (Global Perception $\rightarrow$ Local Scrutiny $\rightarrow$ Critical Thinking) to correctly distinguish artifacts from natural effects. (f-g) Distribution of the Vista-Bench dataset and performance comparison showing Q-Probe’s superiority.
1 Introduction
Image Quality Assessment (IQA) serves as a pivotal fundamental technique in computer vision, designed to emulate human perception of visual fidelity (Fang et al., 2020; Saha et al., 2023; Jia et al., 2025). Within this domain, No-Reference (NR) methods (Wang, 2021; Mao et al., 2025) are highly valued in practical scenarios due to their independence from pristine reference images. However, the transition from traditional handcrafted features (Ahmed et al., 2019) to deep learning paradigms (Talebi and et, 2018; Yu et al., 2024) has long been plagued by challenges regarding overfitting and out-of-distribution generalization (Zhong et al., 2024). The emergence of Multimodal Large Language Models (MLLMs) (Wang et al., 2024, 2025b) has provided a novel solution to this dilemma. Through the Chain-of-Thought (CoT) mechanism, MLLMs integrate low-level distortions (e.g., noise, blur) and high-level semantics (e.g., aesthetics, content), markedly improving model generalization (Wu et al., 2024; You et al., 2025). However, current Supervised Fine-Tuning (SFT) paradigms are still hindered by high annotation expenses and shallow reasoning, failing to fully unleash the inherent cognitive potential of MLLMs.
Recent research paradigms are shifting towards leveraging Reinforcement Learning (RL) to further align models with human perception (Cai et al., 2025; Zhao et al., 2025). This direction advocates for the explicit design of reward functions to unify the complementary dimensions of “response consistency” and “preference alignment,” thereby overcoming the limitations of traditional approaches that optimize solely for ranking or regression accuracy. Two dominant RL paradigms have emerged in IQA. Regression-based methods (e.g., Q-Insight (Li et al., 2025)) treat IQA as absolute score prediction, employing verifiable tolerance rewards to align accuracy with human preference. Conversely, comparison-based methods (e.g., VisualQuality-R1 (Wu et al., 2025)) derive quality from relative differences. By optimizing ranking distributions against human preferences, they induce emergent CoT capabilities that balance ranking accuracy with logical interpretability.
Nevertheless, RL-based IQA methods suffer from a heavy reliance on global macroscopic views, making it difficult to precisely capture subtle local degradation features (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(a)). Even when such artifacts are detected, they are frequently overlooked due to insufficient attentional allocation. In real-world scenarios, degradation in high-value semantic regions (e.g., faces or license plates) should trigger a severe penalty in the quality metric, whereas existing models frequently overestimate quality in such cases due to semantic robustness bias (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(b)). Moreover, these methods are typically restricted to processing global quality. In high-resolution contexts, this coarse-grained approach fails to resolve small objects and subtle local artifacts, leading to marked biases in quality prediction.
The recent emergence of the “Thinking with Images” (OpenAI, 2025) paradigm has spurred the use of multi-scale tools for high-resolution tiny object perception (Zheng et al., 2025; Wang et al., 2025a). While this offers a promising avenue for fine-grained perception in IQA, directly transposing this local zoom-in strategy presents severe challenges. On one hand, traditional visual zoom-in methods relying on region-level supervision benefit from the diversity of target objects, merely requiring SFT trajectories to encompass the target regions. In the IQA context, however, training exclusively on degraded regions is prone to inducing overfitting, causing the model to establish a spurious causal correlation that “cropping implies low quality” (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(d)). Consequently, the model loses its discriminative capability regarding content, exhibiting a systematic low-score bias across all local views. On the other hand, although existing MLLMs excel at local semantic analysis, they often overlook the critical dependency of quality assessment on global context. Taking the Depth-of-Field (DoF) effect in photography as an example, a sharp foreground contrasted with a blurred background typically signifies high-quality imaging. Yet, when isolated from the holistic view and focused solely on the background, models tend to misinterpret natural optical bokeh as degradation (e.g., blur), thereby erroneously assigning a low quality score (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(c)). Furthermore, the current IQA landscape lacks a benchmark dedicated to local fine-grained degradations in high-resolution scenarios, which significantly constrains the evolution of algorithms toward more refined perception.
To address the aforementioned challenges, we propose Q-Probe (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(e)), the first agentic IQA model grounded in the Thinking with Images paradigm and designed specifically for high-resolution scenarios. To enable such fine-grained perception given the lack of dedicated benchmarks, we first construct Vista-Bench (Fig. Q-Probe: Scaling Image Quality Assessment to High Resolution via Context-Aware Agentic Probing(f)), covering a wide range of scenes and local distortion patterns. To ensure the authenticity and precision of the synthesized artifacts, we employ the wavelet transform (Zhang, 2019) to decouple image structure from texture, selectively injecting degradations into texture-rich regions to simulate realistic impairments. We then leverage the advanced capabilities of Gemini-2.5 Pro (Comanici et al., 2025) to achieve hierarchical scores, calculating weighted assessments based on local degradation severity and the semantic significance of the regions. Finally, a rigorous human quality review and filtering process is conducted to guarantee the reliability of the benchmark.
Building on this foundation, we introduce a three-stage training paradigm mimicking the human visual mechanism of “global perception, local scrutiny,” enabling Q-Probe to dynamically optimize tool invocation for comprehensive assessment across resolution scales. Initially, we conduct RL pre-training using a subset of low-resolution images to establish a foundational perception of image quality. This phase is designed to endow the model with the capability to distinguish between actual image degradation and natural DoF bokeh, ensuring that the former is penalized as quality impairment while the latter is recognized as a valid expression of photographic focus.
Following this, we curate Probe-CoT-3K, a dataset comprising mixed-resolution imagery and reasoning trajectories, to facilitate an SFT warm-up phase. Here, we devise a generation strategy that integrates a global overview with diverse local cropping mechanisms (e.g., degradation capture, clarity localization, and distant view assessment). Specifically, the trajectories simulate an initial global reasoning phase, utilizing a holistic view to determine the necessity of further fine-grained perception. During the local cropping phase, unlike traditional VQA methods that focus solely on target semantics, we mandate that cropped regions encompass all degraded patches while simultaneously preserving areas exhibiting pristine natural DoF or clear foregrounds. This design not only effectively eliminates the spurious causal correlation associating “local cropping” with “low quality” but also empowers the model to discern the optimal timing of tool invocation, grounded in a balanced integration of global and local perspectives. However, the strategy of incorporating background context to ensure logical robustness inevitably reduces the model’s precision in localizing subtle degradations, requiring further optimization. To maximize the model’s capacity to identify such fine-grained artifacts, we construct the Probe-RL-4K dataset and implement Reinforcement Fine-Tuning, synchronously optimizing reasoning trajectories and tool invocation precision for accurate decision-making in complex scenarios. Extensive experimental results demonstrate that Q-Probe significantly outperforms existing models in high-resolution settings, accompanied by varying degrees of performance gains in low-resolution scenarios.
Our contributions are summarized as follows:
  • We propose Q-Probe, the first agentic IQA model for high-resolution scenarios, leveraging adaptive cropping to achieve superior capability in assessing both global quality and subtle local degradations.
  • We introduce Vista-Bench, the pioneering benchmark dedicated to high-resolution fine-grained degradation analysis, encompassing a broad spectrum of scenes and local distortion patterns.
  • We construct the Probe-CoT-3K and Probe-RL-4K datasets, leveraging a context-aware cropping strategy to eliminate spurious correlations and generate high-quality CoT trajectories for mastering tool usage.
  • We devise a progressive three-stage training curriculum that first aligns global perception with human preference, then stabilizes logical reasoning via SFT, and finally leverages decoupled reward guidance to recover localization precision.
2 Related Work
2.1 Image Quality Assessment
Image Quality Assessment (IQA) has evolved from early reliance on handcrafted features based on Natural Scene Statistics (NSS) (Mittal et al., 2012a, b) to data-driven deep learning models. While subsequent CNN and Transformer-based approaches achieved performance gains, they often treated IQA as a black-box regression task. The recent advent of Multimodal Large Language Models (MLLMs) introduced a semantic turn, with methods like Q-Align (Wu et al., 2023) and DeQA-Score (You et al., 2025) leveraging foundation models for quality scoring via supervised fine-tuning (SFT). However, these SFT-based paradigms rely heavily on expensive annotations and often suffer from shallow reasoning patterns. To address these limitations, Reinforcement Learning (RL) has emerged as a pivotal mechanism for aligning MLLMs with human preferences. Notably, Q-Insight (Li et al., 2025) pioneered the application of Group Relative Policy Optimization (GRPO) to visual quality understanding, while VisualQuality-R1 (Wu et al., 2025) reformulated IQA as a ranking task via RL-to-Rank.
2.2 Thinking with Images
The “Thinking with Images” paradigm, introduced by OpenAI-o3 (OpenAI, 2025), has spurred the development of agentic visual reasoning MLLMs that interleave image and text reasoning with iterative visual analysis. DeepEyes (Zheng et al., 2025) provides an open-source implementation, demonstrating that end-to-end RL can incentivize models to adopt this behavior for fine-grained tasks. However, subsequent works reveal that pure RL is insufficient for complex, multi-turn interactions. Pixel Reasoner (Wang et al., 2025a) identifies a critical “learning trap” where models bypass nascent visual tools, proposing a two-phase approach: a cold-start phase to establish tool use followed by curiosity-driven RL. Similarly, other works have adopted multi-stage training to activate deep trajectories for hard visual searches. Despite these advances, directly transplanting this agentic probing to IQA remains non-trivial. Unlike object detection, cropping in IQA introduces a causal ambiguity: models trained on cropped degradation patches tend to overfit, learning a spurious correlation that “zooming implies degradation,” thereby penalizing high-quality crops. Q-Probe addresses this gap by introducing a context-aware probing mechanism.
Figure 1: This diagram illustrates the construction pipeline of Vista-Bench and the Data Flywheel for SFT. Specifically, we utilize wavelet transforms to decouple structure from texture, selectively injecting artifacts into texture-rich semantic regions, while employing Gemini-2.5 Pro to generate importance-weighted annotations for fine-grained perception probing. To support SFT, we generate traces that interleave global overviews, defect zooming, and context verification (scrutinizing clear regions), thereby preventing the model from associating tool usage solely with defects.
Figure 2: Overview of the three-stage training framework. Initially, RL Pre-training leverages ranking rewards to align global perception with human preferences. Subsequently, hybrid-resolution SFT enables the model to acquire robust logical reasoning. Finally, the RL Post-training stage fine-tunes the model for precise degradation detection and adaptive tool invocation.
3 Methodology
3.1 Overview
Q-Probe emulates the human “coarse-to-fine” visual mechanism via a progressive three-stage curriculum, balancing global aesthetic perception with local defect scrutiny. To facilitate this, we constructed Vista-Bench, a high-resolution benchmark derived from public datasets through wavelet-based artifact injection and Gemini-2.5 Pro (Comanici et al., 2025) annotation. The training pipeline of Q-probe begins with perception alignment using Group Relative Policy Optimization (GRPO) (Guo et al., 2025) to anchor global judgment. It then progresses to a hybrid-resolution SFT phase driven by a Data Flywheel, employing context-aware cropping to dissociate tool usage from negative quality bias. Finally, a decoupled RL post-training stage utilizes a bifurcated reward mechanism to jointly optimize defect localization and scoring, resolving the conflict between exploration diversity and detection precision.
3.2 Data Construction: Vista-Bench
Existing datasets usually consist of low-resolution images (<1<1K) where degradations are typically applied globally rather than to localized regions. To evaluate fine-grained perception in the high-resolution quality assessment era, we constructed the Vista-Bench.
The dataset is curated from high-quality samples within the HR-Bench 4K (Wang et al., 2025c), Pixel-Reasoner (Wang et al., 2025a) and UHD-IQA (Hosu et al., 2024) datasets. As shown in Fig. 1, the construction pipeline proceeds as follows: we first utilize wavelet transforms to extract global high-frequency components and apply specific degradations (e.g., blur, compression, mosaic). This approach ensures that the degradation is applied to the semantic foreground regions as much as possible. Subsequently, we employ Gemini-2.5 Pro to generate fine-grained quality scores based on the semantic importance of the cropped regions. The resulting benchmark encompasses over 1,000 high-resolution (>4>4K) images, characterized by authoritative, multi-domain, and locally annotated degradations.
3.3 Stage 1: Perception Alignment via Pre-RL
Before introducing complex tool usage, the model must align its fundamental aesthetic perception to overcome “Semantic Robustness Bias.” We employ a dataset of 3,000 low-resolution images randomly sampled from the KADID-10k (Lin et al., 2019) dataset to train the model via a pairwise variant of GRPO.
Probabilistic Ranking via Thurstone Model. Unlike standard regression approaches that treat quality scores as deterministic point estimates, we model visual quality as an intrinsically relative concept governed by the Thurstone model (Thurstone, 2017) case V. For a training pair of images $\{x_{i},x_{j}\}$ (effectively a batch size of $B=2$), the policy model $\pi_{\theta}$ generates $K$ distinct reasoning paths and quality scores for each image: $q(x_{i})=[q_{1}(x_{i}),\dots,q_{K}(x_{i})]^{\intercal}$. This distribution inherently captures predictive uncertainty, which serves as a cornerstone for reliable ranking.
We calculate the asymmetric comparative probability that $x_{i}$ is perceptually superior to $x_{j}$ by standardizing the difference in their predicted means against their joint uncertainty. To improve numerical stability, we first define the standardized difference score $\mathcal{Z}_{ij}^{k}$:
$\mathcal{Z}_{ij}^{k}=\frac{q_{k}(x_{i})-\mu(q(x_{j}))}{\sqrt{\sigma^{2}(q(x_{i}))+\sigma^{2}(q(x_{j}))+\gamma}}$ (1)
where $\mu(\cdot)$ and $\sigma^{2}(\cdot)$ denote the sample mean and variance of the predicted scores group, $q_{k}(x_{i})$ represents the $k$-th score estimate for image $x_{i}$, and $\gamma$ is a small constant ensuring numerical stability. The comparative probability is then derived via the standard Gaussian Cumulative Distribution Function (CDF):
$P_{k}(x_{i}\succ x_{j})=\Phi\left(\mathcal{Z}_{ij}^{k}\right)$ (2)
where $\Phi(\cdot)$ is the Gaussian CDF, representing the probability $P_{k}$ that image $x_{i}$ is perceptually superior to $x_{j}$. Crucially, we explicitly leverage the sample variances derived from the GRPO group outputs to dynamically accommodate predictive uncertainty, preventing over-confident errors on ambiguous images.
Fidelity Reward and Optimization. To rapidly align the model’s predictions with human perception, we define the ground truth preference $y_{ij}$ based on Mean Opinion Scores (MOS): $y_{ij}=1$ if $\text{MOS}(x_{i})>\text{MOS}(x_{j})$, $0.5$ if equal, and 0 otherwise. The optimization is driven by a continuous fidelity reward. We define the ranking reward $R_{rank}$ as:
$\begin{split}R_{rank}(x_{i})=\ &y_{ij}\cdot P_{k}(x_{i}\succ x_{j})\\ &+(1-y_{ij})\cdot(1-P_{k}(x_{i}\succ x_{j}))\end{split}$ (3)
where $y_{ij}\in\{0,0.5,1\}$ represents the ground truth preference label derived from human ratings. This reward is designed to capture fine-grained distinctions in quality ranking. To ensure training stability, we minimize the GRPO objective $\mathcal{J}(\theta)$, which integrates the clipped surrogate advantage and a KL-divergence penalty:
$\mathcal{J}(\theta)$ $=-\mathbb{E}\Bigg[\min\left(\frac{\pi_{\theta}}{\pi_{old}}A_{k},\text{clip}\left(\frac{\pi_{\theta}}{\pi_{old}},1-\epsilon,1+\epsilon\right)A_{k}\right)$
$\qquad-\beta D_{KL}(\pi_{\theta}||\pi_{ref})\Bigg]$ (4)
where $\pi_{\theta}$ and $\pi_{old}$ represent the current and previous policies respectively, $A_{k}$ is the advantage score computed from reward signals, $\epsilon$ is the clipping threshold to constrain policy updates, $\pi_{ref}$ is the reference policy (typically the initial SFT model), and $\beta$ is a coefficient controlling the strength of the KL-divergence penalty $D_{KL}$. This pre-alignment stage establishes a solid baseline for global image understanding, converging towards the mechanisms of human MOS.
3.4 Stage 2: Hybrid-Resolution SFT via Data Flywheel
To bridge the gap between global perception and local scrutiny, we construct the Probe-CoT-3K dataset using a Data Flywheel approach (as illustrated in Fig. 1). The dataset follows a 2:1 ratio of low-to-high resolution images.
The SFT Dilemma: Precision vs. Diversity. In traditional Visual Question Answering (VQA), crop content inherently exhibits high diversity. However, in Agentic IQA, if SFT data is invariably centered on degradations, the model overfits to a biased mode: “Tool Usage implies Degradation and Low Quality,” resulting in a fundamental breakdown of logical reasoning.
Context-Aware Trajectory Generation. To mitigate this, we synthesize CoT trajectories encompassing genuine degradations, natural depth-of-field, and pristine foregrounds. This strategy effectively severs the spurious causal correlation associating “cropping” with “low quality”. We optimize the model using the standard cross-entropy loss over the mixed sequence of reasoning and action tokens $y$:
$\mathcal{L}_{SFT}=-\sum_{t=1}^{L}\log P(y_{t}|y_{<t},\mathbf{x})$ (5)
where $\mathcal{L}_{SFT}$ is the SFT loss, $L$ denotes the sequence length, $\mathbf{x}$ represents the multimodal input (image and instruction), and $y_{t}$ is the $t$-th token in the generated target sequence $y$ given the preceding tokens $y_{<t}$. Following the SFT phase, the model establishes a robust understanding of the global image context and acquires the capability to accurately discern diverse local regional features.
3.5 Stage 3: Precision Pursuit via Decoupled Post-RL
As the SFT phase strategically trades localization precision for logical robustness, we construct and leverage the Probe-RL-4K dataset in the final stage to further refine the model’s fine-grained detection capabilities.
Decoupled Reward Mechanism. We implement a decoupled reward mechanism that disentangles the optimization of the “Looking Policy” from the “Scoring Policy.” We independently define the accuracy reward $R_{acc}$ and the localization reward $R_{loc}$:
$R_{acc}$ $=\exp\left(-\frac{|s_{pred}-s_{MOS}|}{\tau}\right)$ (6)
$R_{loc}$ $=\mathbb{I}(has\_defect)\cdot\text{IoU}(B_{pred},B_{gt})$ (7)
where $s_{pred}$ and $s_{MOS}$ denote the predicted score and ground truth MOS respectively, and $\tau$ is a temperature hyperparameter controlling the sensitivity of the accuracy reward. For localization, $\mathbb{I}(\cdot)$ is an indicator function that equals 1 if a defect exists in the image, and $\text{IoU}(\cdot)$ computes the Intersection over Union between the predicted crop box $B_{pred}$ and the ground truth defect region $B_{gt}$. This reward incentivizes the model to accurately capture degradations and precisely deploy tools for regional magnification.
The total reward $R_{total}$ is then composed as:
$R_{total}=\underbrace{\alpha R_{acc}}_{\text{Scoring}}+\underbrace{\beta R_{loc}}_{\text{Looking}}+\gamma R_{format}$ (8)
where $\alpha$, $\beta$, and $\gamma$ are hyperparameters weighting the contribution of the accuracy reward, localization reward, and format compliance reward $R_{format}$, respectively. This decoupled reward optimizes the model to refine its localization capabilities ($\beta R_{loc}$), building upon the robust logic reasoning capability acquired during SFT.
Table 1: Comprehensive performance comparison (SRCC / PLCC) on standard and high-resolution datasets. Q-Probe achieves SOTA results, especially on the high-resolution Vista benchmark. Methods are color-coded by category: Handcrafted, Deep Learning, and MLLMs-based. The best and second-best results are highlighted in red and blue.
Methods Vista SPAQ KADID PIPAL TID13 KonIQ AGIQA Avg
Spearman Rank Correlation Coefficient (SRCC)
[…]
Pearson Linear Correlation Coefficient (PLCC)
[…]
Figure 3: To monitor the training dynamics, we calculated the average standard deviation of predicted scores across multiple inference runs at various checkpoints. The observed monotonic decrease in variance not only confirms that Q-Probe achieves greater stability throughout Stage-1.
Table 2: Ablation study of the progressive training strategy across multiple benchmarks.
Model Configuration Vista SPAQ KADID-10k KonIQ-10k Average
SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Stage 1 Only 0.290 0.340 0.760 0.780 0.700 0.725 0.650 0.680 0.600 0.631
Stage 2 + Stage 3 0.679 0.722 0.766 0.796 0.781 0.772 0.752 0.769 0.746 0.755
Stage 1 + Stage 2 0.690 0.735 0.850 0.860 0.880 0.875 0.820 0.815 0.810 0.821
Full (Stage 1+2+3) 0.728 0.776 0.892 0.900 0.901 0.892 0.871 0.863 0.848 0.858
Table 3: Ablation of the Reward Mechanism in Stage 3 on Vista-Bench.
Reward Strategy SRCC PLCC
$R_{acc}$ (Score Only) 0.698 0.747
$R_{acc}+R_{format}$ 0.705 0.752
$R_{acc}+R_{format}+R_{loc}$ (Ours) 0.728 0.776
Table 4: Ablation of Crop Coverage Strategies regarding Degradation Regions on Vista-Bench.
Crop Strategy SRCC PLCC
Degradation Only 0.510 0.580
Partial Degradation & Partial Normal 0.695 0.745
All Degradation & Partial Normal 0.728 0.776
4 Experiments
4.1 Experimental Setup
Datasets and Metrics. To evaluate the comprehensive capabilities of Q-Probe, we conduct experiments across a diverse set of IQA benchmarks. For standard resolution assessments, we utilize KonIQ-10k (Hosu et al., 2020), SPAQ (Fang et al., 2020), and KADID-10k (Lin et al., 2019) to represent in-the-wild and synthetic distortions. We also include PIPAL (Jinjin et al., 2020) to evaluate robustness against algorithm-processed distortions and AGIQA-3K (Li et al., 2023) for AI-generated content. Crucially, to validate fine-grained perception in the high-resolution era, we employ our proposed Vista-Bench, which contains localized artifacts in high-resolution scenarios. We leverage the widely recognized Spearman Rank-order Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC) as our primary evaluation metrics. Following established protocols (Wu et al., 2023; Fang et al., 2020), we compute the average performance across datasets to assess overall generalization.
Baseline Methods. We compare the proposed Q-Probe against a wide range of state-of-the-art (SOTA) baseline methods, including handcrafted methods (NIQE (Mittal et al., 2012b), BRISQUE (Mittal et al., 2012a)), discriminative deep learning models (UNIQUE (Zhang et al., 2021), MUSIQ (Ke et al., 2021), MANIQA (Yang et al., 2022)), MLLMs-based approaches (LIQE (Zhang et al., 2023), Q-Align (Wu et al., 2023), DeQA-Score (You et al., 2025), and the recent RL-based models Q-Insight (Li et al., 2025) and VisualQuality-R1 (Wu et al., 2025)).
Implementation Details. We utilize Qwen-2.5-VL-7B (Bai et al., 2025) as our base model. During the perception alignment stage, we employ GRPO with a generation number $K=6$ and a clip threshold $\epsilon=0.2$. The model is trained using the AdamW optimizer with a learning rate of $1\times 10^{-6}$. For the hybrid-resolution SFT, we utilize the Probe-CoT-3K dataset with a crop resolution of $768\times 768$. In the decoupled Post-RL stage, we set the reward weights $\alpha=1.0$ (accuracy) and $\beta=0.15$ (localization). All experiments are conducted on 8 NVIDIA A100 (80GB) GPUs.
4.2 Main Results
Table 1 presents the comprehensive SRCC and PLCC performance across seven benchmarks. In terms of SRCC, our method attains a remarkable score of 0.728 on the challenging Vista-Bench dataset, significantly outperforming the baseline methods which lack fine-grained perception mechanisms. The PLCC results in the lower half of Table 1 confirm the linear alignment of our predictions with human perception, achieving a leading PLCC of 0.776 on high-resolution scenarios. Moreover, experimental results demonstrate that our method maintains superior performance across diverse low-resolution datasets, underscoring the robust generalization capability of Q-Probe across resolution scales.
4.3 Ablation Studies
Necessity of the Three-Stage Curriculum.
Table 2 validates the efficacy of our progressive training curriculum. Although Stage 1 performs poorly on high-resolution details, it successfully pre-aligns the model with human aesthetic perception on standard benchmarks. Without Stage 1, model can’t reach its best performance. Building on this foundation, Stage 2 bridges the granularity gap via hybrid-resolution training, while Stage 3 further refines the scoring precision through RL to achieve the best overall performance. Besides, three Stages consistently boost performance on standard low-resolution benchmarks, culminating in the best overall results.
Impact of Reward Components in Post-RL.
In Stage 3, we analyze the contribution of different reward components. As shown in Table 3, using only the Accuracy Reward ($R_{acc}$) provides a solid baseline. Incorporating the Format Reward ($R_{format}$) yields a slight improvement by ensuring the structural validity of the CoT reasoning. However, the most significant gain comes from the Decoupled Localization Reward ($R_{loc}$), which explicitly guides the model’s attention to defects, pushing the SRCC to 0.728.
Impact of Crop Coverage Strategy.
Table 4 examines the spatial relationship between crops and degradation. Restricting crops to exact degradation regions (Strategy 1) leads to severe overfitting, where the model learns biased inference and associates all crops with degration, dropping SRCC to 0.510. Strategy 2, covering partial degradation and normal areas, improves to 0.695 but introduces trade-off that the model may overlook omitted defects. Finally, Strategy 3 achieves optimal performance (SRCC 0.728) by capturing all degradation within a normal context, ensuring comprehensive defect assessment without hallucinations.
5 Conclusion
In this work, we introduced Q-Probe, the first agentic framework designed to scale IQA to high-resolution scenarios via context-aware probing. Recognizing the limitations of existing global-view RL methods in capturing fine-grained artifacts, we proposed a novel three-stage training curriculum that mimics the human “coarse-to-fine” perception mechanism, teaching the model to evaluate quality based on importance of degraded crops. Furthermore, by constructing the Vista-Bench and leveraging a Data Flywheel, we eliminated the spurious “cropping-implies-degradation” bias, enabling the model to intelligently distinguish between technical distortions and artistic effects. Our model achieves sota performance on both high-resolution and low-resolution datasets.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Data Construction and Automated Annotation Details
To efficiently construct Vista-Bench and scale Image Quality Assessment (IQA) to high-resolution scenarios, we developed a high-throughput, asynchronous data processing pipeline. This system leverages Gemini-2.5-Pro to generate fine-grained, hierarchical quality annotations. This section details the processing workflow and the specific prompt engineering strategies employed.
A.1 Wavelet-Based Degradation Injection
Existing IQA datasets often rely on global distortions (e.g., Gaussian blur applied uniformly), which allow models to bypass genuine visual scrutiny by learning simple global statistical deviations. To prevent such “shortcut learning” and simulate realistic high-resolution defects, we employ a targeted degradation injection strategy based on the Discrete Wavelet Transform (DWT).
Structure-Texture Decoupling.
We utilize the Haar wavelet to decompose high-resolution source images into four distinct frequency subbands: the low-frequency approximation ($LL$) and three high-frequency detail components ($LH$, HLHL, HHHH). The $LL$ band preserves the global semantic structure and flat regions, while the high-frequency bands capture fine-grained textures and edges where high-resolution artifacts typically manifest, as shown in Figure 4.
Targeted Artifact Synthesis.
Instead of corrupting the entire image, we selectively inject degradations (e.g., compression ringing, sensor noise, and slight defocus) specifically into the high-frequency subbands. This strategy ensures that artifacts are embedded within complex texture regions—such as hair, foliage, or architectural details—making them visually subtle yet perceptually critical. By keeping the $LL$ band intact, we maintain the overall structural integrity of the scene while introducing realistic local impairments.
Reconstruction.
The final degraded samples are generated via the Inverse Discrete Wavelet Transform (IDWT). This process forces the IQA agent to actively “zoom in” and scrutinize local details to detect quality drops, effectively simulating the fine-grained perception required for 4K scenarios.
A.2 Asynchronous Annotation Pipeline
Given the high-resolution nature of the source images and the need for dense cropping, a sequential processing approach would be computationally prohibitive. We implemented a Python-based asynchronous framework tailored for high-concurrency interaction with the Multimodal LLM (MLLM). The pipeline consists of three critical stages:
  1. 1.
    Dynamic Image Preprocessing and Encoding: To optimize the balance between visual fidelity and token consumption, we implemented an adaptive resizing strategy. Original high-resolution images are resized such that their maximum edge does not exceed 1,024 pixels ($H_{max}=1024$) using Lanczos resampling. This ensures that global artifacts are preserved while significantly reducing bandwidth usage. Images are then serialized into Base64 format for API transmission.
  2. 2.
    Semaphore-Based Concurrency Control: To maximize throughput without exceeding API rate limits, the system utilizes an asynchronous semaphore mechanism (set to a concurrency limit of 5). This allows for parallel processing of image batches while maintaining system stability. The pipeline utilizes aiohttp for non-blocking network I/O, ensuring that image encoding and API waiting times do not bottle-neck the annotation process.
  3. 3.
    Robust Output Parsing: Since MLLM outputs can vary in formatting, we implemented a resilient parsing logic that utilizes regular expressions to extract JSON structures from mixed-text responses. This ensures valid data extraction even if the model wraps outputs in Markdown fences or custom XML-like tags.
A.3 Context-Aware Prompt Design
A critical challenge in Agentic IQA is preventing the model from hallucinating artifacts in clean, high-quality crops (e.g., natural bokeh). To address this, we designed a Two-Step Context-Aware Prompt that enforces a “Zero-Penalty” rule.
The prompt explicitly instructs the agent to first establish a global baseline score and then apply penalties only if specific, man-made degradations are detected. The structured system instructions provided to the agent are presented below.
Core Prompt for Context-Aware Annotation SYSTEM INSTRUCTION:
You are an advanced Image Quality Assessment Expert.
USER INSTRUCTION:
You are provided with multiple images.
Image 1 is the FULL ORIGINAL IMAGE. The remaining images are CROPPED VIEWS (zoomed-in details) from Image 1. CRITICAL KNOWLEDGE: TARGET DEGRADATIONS
You are specifically looking for 5 types of MAN-MADE degradations:
1. Mosaic: Pixelation blocks. 2. JPEG: Heavy compression ringing/blocks. 3. Posterize: Banding or flat color areas. 4. Scanlines: Horizontal black lines. 5. Grid: White/colored grid lines overlaid. All these 5 degradations must be artificial blocks, not natural textures. Step 1: Analyze Global Quality (Baseline)
Evaluate global tech-fidelity. Do NOT penalize natural bokeh/blur. Start with a Base Score (1.0 - 5.0). High-quality usually starts at 4.5+.
Step 2: Crop Analysis & “Zero-Penalty” Rule
Examine each crop carefully.
If a crop is Clean (no artificial degradation): Penalty = $0.$ If a crop is Degraded: Apply penalty based on location and severity. Output Format
Please strictly follow this JSON structure:
  "step1_base_score": 4.8,
  "step2_crop_details": [
      "crop_id": "Crop 1",
      "status": "Clean",
      "reasoning": "Natural leaf texture...",
      "penalty": 0.0
      "crop_id": "Crop 2",
      "status": "Degraded",
      "reasoning": "Distinct 8x8 mosaic blocks...",
      "penalty": 0.5
  "step3_summary": "Base score 4.8 minus penalties...",
  "final_score": 4.3
This hierarchical prompting strategy ensures that the final score (final_score) is mathematically derived from the baseline (step1_base_score) minus the localized penalties, effectively decoupling artistic blur from technical degradation.
Appendix B Inference Pipeline Implementation Details
To evaluate Q-Probe’s capability to detect fine-grained degradations in high-resolution images, we developed a specialized inference pipeline based on the vLLM framework. This pipeline supports multi-turn visual reasoning and dynamic tool invocation, allowing the model to actively ”zoom in” on suspicious regions. The implementation details are described below.
Figure 4: We utilize the Wavelet Transform to extract high-frequency components, serving as spatial guidance to concentrate the degradation injection within semantically rich foreground regions.
B.1 Engine Configuration and Resolution Management
We deployed the model using the vLLM engine to maximize inference throughput and memory efficiency. To accommodate the ultra-high resolution of Vista-Bench images (often exceeding 4K resolution), we customized the multi-modal processor configuration. Specifically, we set the min_pixels to $256\times 28\times 28$ and max_pixels to $2560\times 28\times 28$. This configuration ensures that the model preserves sufficient high-frequency details during the initial global encoding phase, preventing the loss of small artifacts (e.g., subtle mosaic blocks or grid lines). The inference was conducted using bfloat16 precision with a tensor parallel size of 2, distributed across NVIDIA RTX 4090 GPUs. We also enabled enable_prefix_caching to accelerate the processing of shared system prompts across multiple turns.
B.2 The Agentic Reasoning Loop
Unlike traditional static inference, our pipeline implements a dynamic, multi-turn ”Observe-Reason-Act” loop. The process is defined as follows:
  1. 1.
    Initialization: The session begins with the system prompt defining the five target degradation types (Mosaic, JPEG, Posterize, Scanlines, Grid) and the ”Zero-Penalty” rule. The global view of the high-resolution image is fed into the model.
  2. 2.
    Global Analysis and Hypothesis Generation: The model first evaluates the global technical fidelity. If it detects potential local degradations that are ambiguous in the global view, it generates a special XML tag <tool_call> containing the normalized bounding box coordinates for a closer inspection.
  3. 3.
    Dynamic Tool Execution: A regular expression parser monitors the model’s output stream. Upon detecting a crop_image_normalized function call, the pipeline pauses text generation and executes the cropping logic:
    $I_{crop}=\text{Crop}(I_{global},[x_{1},y_{1},x_{2},y_{2}])$ (9)
    The cropped region is extracted from the original high-resolution source (not the resized tokens) to ensure maximum clarity.
  4. 4.
    Contextual Update: The cropped image is encoded into Base64 format and appended to the conversation history as a new user message. The model then resumes generation, analyzing the new visual evidence to confirm or dismiss the degradation hypothesis.
  5. 5.
    Termination: This loop continues for a maximum of $T=6$ turns. The process terminates early if the model ceases to call tools or provides a final score.
B.3 State Management and Reproducibility
[…]
B.4 Prompt Engineering
To guide the agent in performing rigorous technical quality assessment, we designed a structured System Prompt. This prompt serves four key functions: Role Assignment, Defect Taxonomy, Workflow Enforcement, and Output Standardization.
The exact system prompt used in our inference pipeline is shown below.
System Prompt for Agentic IQA Inference You are an expert Image Quality Assessment (IQA) assistant. Your task is to analyze the image for technical degradations, specifically focusing on digital artifacts like: 1. Mosaic/Pixelation: Blocky structures. 2. JPEG Compression: Ringing artifacts and blocking. 3. Posterization: Color banding or loss of gradient. 4. Scanlines/Grid: Artificial line overlays. 5. Noise/Blur: General degradation (common in KADID10K). Inference Steps: 1. Global Analysis: Briefly describe the overall content and perceived quality. Is it sharp? Is it noisy? 2. Defect Hunting: Actively look for the artifacts listed above. 3. Zoom/Crop Strategy: If you suspect an artifact in a specific area (e.g., background textures, edges), use the crop_image_normalized tool to inspect it closely. 4. Final Scoring: Based on the severity of the defects found: 5.0: Excellent, no visible defects. 4.0: Good, minor defects visible only upon close inspection. 3.0: Fair, defects are annoying but content is visible. 2.0: Poor, defects significantly affect quality. 1.0: Bad, extremely degraded. CRITICAL OUTPUT RULE: At the very end of your final response, you MUST output the final score enclosed in double brackets, strictly like this: [[3.45]].
Appendix C Supervised Fine-Tuning Data Construction
To equip the Q-Probe agent with both robust visual perception and reasoning capabilities, we constructed a high-quality Supervised Fine-Tuning (SFT) dataset, named Probe-CoT-3K. This dataset is designed to teach the model not only to detect degradations but also to articulate its reasoning process in a structured manner. The construction process involves a rigorous template-based generation strategy that pairs diverse images with multi-turn conversation trajectories.
C.1 Data Format and Templating
Each data entry in Probe-CoT-3K follows a standardized multi-turn conversation format, simulating the interaction between a user and an expert IQA agent. The data structure is defined as follows:
  • System Prompt: A fixed instruction set that defines the agent’s persona, the specific degradation taxonomy (Mosaic, JPEG, Posterize, Scanlines, Grid), and the critical “Zero-Penalty” rule for clean crops.
  • Global Analysis Turn: The first turn involves the user presenting the full image. The assistant’s response includes a global quality assessment, a tentative base score, and—crucially—the invocation of the crop_image_normalized tool to inspect suspicious regions.
  • Local Scrutiny Turn: The subsequent turn simulates the system returning the cropped images. The assistant then analyzes these local views, identifies specific artifacts (if any), applies the appropriate penalties, and calculates the final quality score.
C.2 Trajectory Diversity and Quality Control
A key challenge in SFT for agentic tasks is avoiding “shortcut learning,” where the model blindly predicts low scores whenever a crop tool is used. To mitigate this, our dataset includes a balanced mix of three trajectory types:
  1. 1.
    Positive Trajectories: The agent zooms in on complex textures (e.g., foliage, fabric) and correctly identifies them as clean, applying zero penalty.
  2. 2.
    Negative Trajectories: The agent detects actual degradations (e.g., JPEG blocks) in the crops and applies penalties.
  3. 3.
    Global-Only Trajectories: For obviously low-resolution or low-quality images, the agent decides not to zoom in and provides a score directly based on the global view.
Table C.2 illustrates the template structures used for generating these reasoning trajectories, showcasing the distinction between the agentic zoom-in path and the direct scoring path.
Table 5: The template structures for the SFT data. To ensure model robustness, we construct two distinct reasoning paths: (a) The Agentic Zoom-in Path for high-resolution images requiring local scrutiny, and (b) The Direct Scoring Path for low-resolution.
(a) High-Res / Agentic Path (With Crop Tool Usage)
[…]
Appendix D More Details about Vista-Bench
The significant performance superiority of Q-Probe over existing global-view IQA methods on Vista-Bench can be attributed to the fundamental paradigm shift from passive observation to active Context-Aware Agentic Probing. While traditional methods rely on coarse-grained global views that often miss high-frequency details, our framework introduces a nuanced evaluation capability centered on two key advantages:
1. Fine-Grained Perception vs. Coarse Global Scoring: Existing MLLMs typically process images from global perspective, causing subtle degradations in high-resolution scenarios (e.g., $4K$ inputs) to vanish during encoding. These models effectively ”guess” the quality based on a low-fidelity global impression. In contrast, Q-Probe utilizes an agentic zooming mechanism to actively localize and inspect microscopic degradations. Crucially, our method does not merely detect defects; it evaluates the perceptual impact of the degradation relative to the semantic importance of the region. By integrating the severity of local defects with the saliency of the content (e.g., applying severe penalties for artifacts on a foreground subject while adhering to a ”Zero-Penalty” rule for natural background bokeh), Q-Probe aligns its scoring logic with human visual attention, a capability largely absent in global-view baselines.
2. Generalization to Unseen Degradations: To rigorously validate that the performance gains stem from genuine visual understanding rather than distribution overfitting, we strategically incorporated unseen degradation types into the Vista-Bench test set—distortion patterns that the model was never exposed to during training. Standard baselines often fail on such out-of-distribution data as they rely on memorizing specific artifact features. However, Q-Probe demonstrates sustained high performance on these unseen samples. This proves that our model has transcended simple pattern matching and has acquired a robust, abstract capability to assess how generic visual anomalies impair the overall integrity of an image, regardless of whether the specific degradation pattern has been previously encountered.
Appendix E Limitations and Bad Cases Study
To comprehensively evaluate the boundaries of Q-Probe, we analyze both successful applications and failure modes. As demonstrated in Figure 5, under standard high-resolution scenarios, our agent exhibits superior performance. The context-aware cropping strategy successfully distinguishes between artificial degradations and natural depth-of-field effects (bokeh). By accurately localizing low-quality regions while ignoring non-subject blurry backgrounds, the model adheres strictly to the ”Zero-Penalty” rule, ensuring fair and accurate quality assessment.
Figure 5: Success Case: The agent effectively identifies the main subject and applies the “Zero-Penalty” rule to the natural bokeh in the background, demonstrating the effectiveness of our context-aware cropping strategy in standard scenarios.
However, robustness challenges arise in scenes exhibiting complex semantic dependencies. The first limitation involves semantic entanglement between foreground and background. As illustrated in Figure 6, while the model ignores standalone blurry backgrounds, it struggles when distant backgrounds contain textures strongly correlated with the subject’s narrative (e.g., a complex street scene behind a vehicle).
In these cases, the agent’s attention mechanism becomes confused by the high correlation. Instead of focusing solely on the subject, it erroneously interprets complex background textures as ”artificial noise” or ”processing artifacts.” This leads to erroneous cropping on irrelevant areas and subsequently applies unjustified penalties, indicating a need for a more sophisticated semantic understanding of scene composition.
Figure 6: Failure Mode I (Semantic Entanglement): While the agent can distinguish bokeh, its discriminative capability weakens when the background is strongly correlated with the subject (e.g., a car on a street). The model may misinterpret complex background textures as artifacts, leading to erroneous cropping.
The second limitation pertains to the insensitivity to spatial context differences, particularly between indoor and outdoor environments. As shown in Figure 7, our agent is prone to uniformly assigning minimal penalties to degradations in all background regions. This generalization is flawed because indoor backgrounds cannot be treated equivalently to vast outdoor settings.
Due to the confined spatial nature of indoor scenes, background artifacts remain perceptually significant compared to distant outdoor backgrounds. Consequently, while the penalty for indoor background degradation should not be as severe as foreground defects, it must not be negligible either. Our current model fails to apply this necessary moderate penalty, often underestimating the impact of background noise in spatially constrained environments.
Figure 7: Failure Mode II (Spatial Context Insensitivity): The agent incorrectly applies minimal penalties to indoor background degradations. Unlike vast outdoor settings, indoor scenes are spatially confined, meaning background artifacts remain perceptually significant and require a moderate penalty rather than a negligible one.
Finally, a third limitation arises from ambiguity in subject definition, leading to a salience bias during crop selection. As depicted in Figure 8, the agent incorrectly identifies the distant house as the primary subject—likely due to its status as the sole distinct architectural element—while relegating the river to the background.
This classification ignores the fact that the river occupies nearly half the frame and dominates the visual experience. In the context of Image Quality Assessment (IQA), degradations on such a prominent surface significantly impair perceptual quality. By classifying the river as a non-salient background, the model generates a flawed reasoning trajectory and fails to penalize visible defects, resulting in an inflated quality score.
Figure 8: Failure Mode III (Salience Bias): Ambiguity in scene composition introduces a bias: the agent misidentifies the small distant house as the main subject due to object uniqueness, while dismissing the dominant river as background. Consequently, the model fails to penalize degradations on the water surface, which are actually critical to perceptual quality.